OpenStudy (anonymous):

Which expression is equivalent to the area of metal sheet required to make this square-shaped traffic sign? A square shaped traffic sign is shown with length of one side labeled as x + 1. x2 + 2x + 1 x2 + x + 1 x^2 + 2x x^2 + 1

OpenStudy (agent0smith):

Answered this in the last one: If the side length is x+1, then the area is the (side length)^2, or (x+1)^2. (x+1)^2 = (x+1)(x+1) =